# Cache Me If You Must: Adaptive Key-Value Quantization for Large Language Models

### v2

@@ 4 Experiments
When evaluating perplexity, we adopt the same approach as in prior quantization works (Frantar et al., 2022; Lin et al., 2023; Egiazarian et al., 2024), with one exception: instead of processing sequences (of length 8192) in parallel, we encode them auto-regressively and maintain the (compressed) Key-Value cache as during as during inference. This results in the same perplexity for non-compressed KV cache, but allows us to properly account for the effect of recent token buffers and attention sinks during KV quantization, as described in Section 3.3. Here, we use base (non-instruct) models since they have better perplexity.
In turn, LongBench v1 (Bai et al., 2023) contains long-context length evaluation benchmarks including QA tasks, summarization, and few-shot learning. We evaluate all 14 English-language tasks, without restricting input length to 8192 tokens. This allows us to better explore the effectiveness of AQUA-KV on longer sequences. Since 14 individual tasks are often difficult to analyze, we report the average score across all tasks and provide detailed per-task results in Appendices DE & F. We use the official benchmark code, and evaluate on Instruct models since many LongBench tasks were designed for such models (see Appendix C).
@@ 4.1 Detailed Evaluation & Ablation Analysis
Alternative quantizers. As discussed, AQUA-KV is compatible with any “backbone” quantization scheme. We explore two such algorithms: HIGGS and Quanto, as described in Section 3.3. For HIGGS quantization, we use quantization group size 1024 and employ the grid configurations that have fast GPU support for encoding and decoding, at 2-, 3- and 4-bit precision. For Quanto, we use the default group size 64 and per-token compression (0-th axis). Both methods use Round-To-Nearest (RTN) quantization. We compare each algorithm with their respective variant without learned predictors, as well as two popular algorithms for key-value cache compression: KIVI (Liu et al., 2024c) and KVQuant (Hooper et al., 2024).
@@
Uncompressed 16 40 6.98 5.61 2.54 7.14 6.13 44.61 48.13 52.92 31.61 39.14
AQUA-KV 2.09 5.7 7.03 5.72 2.62 7.20 6.17 44.30 47.77 52.79 29.92 39.27
HIGGS 2.02 5.1 7.47 5.89 2.77 7.93 8.08 42.80 47.37 52.18 25.41 19.85
@@ Appendix C On LongBench Evaluation on non-Instruct models
Quanto 2-bit gs64 2.25 21.56 21.14 27.03 12.83 64.00 13.05 10.57 81.21 15.54 0.12 22.30 6.01 11.63 5.51 10.84 15.30
AQUA-KV (Quanto gs64) 2.39 10.33 26.24 41.59 11.36 69.00 10.60 21.85 88.33 22.91 0.00 32.67 6.93 12.97 6.65 12.75 29.69
@@
We provide scores on each of the LongBench tasks in Table 5, as well as some additional experiments. The table is supplemented with results on Quanto with larger group size; on Quanto, applied after RoPE; on KVSharer with 2 shared pairs; on MLP predictors without Layer Nomalization, on keeping longer sequance of unquantized tokens in the beginning of the sentence; on different predictor input configuration.
@@
AQUA-KV (Quanto gs64) 2.39 10.33 43.64 41.29 39.99 71.50 53.72 26.15 88.68 23.86 3.53 50.61 24.96 40.38 88.50 24.26 33.54
AQUA-KV (Quanto gs1024) 2.16 53.35 43.66 42.11 39.53 71.00 51.26 25.84 88.67 24.18 4.06 49.36 26.86 39.52 89.50 25.36 33.96
@@
Before RoPE Quanto gs64 2.39 10.90 44.54 42.57 41.31 71.50 53.10 26.02 88.16 24.01 3.53 50.59 26.91 41.50 95.50 25.10 33.69
After RoPE Quanto gs64 2.39 10.33 43.64 41.29 39.99 71.50 53.72 26.15 88.68 23.86 3.53 50.61 24.96 40.38 88.50 24.26 33.54
@@
Uncompressed 16 7.14 31.61 44.45 20.46 68.00 24.51 23.02 87.60 22.01 0.50 36.46 11.55 20.96 43.00 13.07 26.97
AQUA-KV 2.44 7.20 31.93 43.93 22.02 68.00 27.85 22.30 88.01 21.68 0.50 38.27 12.46 20.97 42.12 13.72 25.23
HIGGS 2.06 7.92 25.41 38.91 14.55 49.00 18.98 19.12 85.94 18.60 0.77 33.34 6.64 17.88 19.67 11.13 21.21
@@
Uncompressed 16 6.13 39.14 45.77 24.29 72.00 29.17 23.53 89.42 21.96 8.00 42.14 14.15 25.34 100.0 20.72 31.42
AQUA-KV 2.48 6.17 39.27 45.99 21.68 72.00 31.36 22.71 89.18 22.03 8.50 43.25 16.92 25.25 99.00 21.18 30.68
HIGGS 2.03 8.08 19.85 26.98 7.86 52.50 10.41 11.18 55.46 12.27 5.58 19.10 4.16 12.50 42.54 5.81 11.61
@@
AQUA-KV 4.43 6.14 39.24 45.81 24.36 72.00 28.98 24.01 89.83 22.30 9.25 42.61 13.98 25.62 99.50 20.15 30.91
HIGGS 4.03 6.88 9.97 11.80 4.68 37.75 5.32 4.57 22.54 10.32 4.06 7.80 2.91 3.83 14.25 5.02 4.72
KIVI 4.25 6.14 33.40 46.2 9.61 71.00 10.34 22.27 89.63 21.13 4.58 30.56 6.85 12.79 98.25 12.06 32.39

### v4

@@ 4 Experiments
When evaluating perplexity, we adopt the same approach as in prior quantization works (Frantar et al., 2022; Lin et al., 2023; Egiazarian et al., 2024), with one exception: instead of processing sequences (of length 8192) in parallel, we encode them auto-regressively and maintain the (compressed) Key-Value cache during inference. This results in the same perplexity for non-compressed KV cache, but allows us to properly account for the effect of recent token buffers and attention sinks during KV quantization, as described in Section 3.3. Here, we use base (non-instruct) models since they have better perplexity.
In turn, LongBench v1 (Bai et al., 2023) contains long-context length evaluation benchmarks including QA tasks, summarization, and few-shot learning. We evaluated all the 14 English-language tasks without restricting the input length to 8192 tokens. This allows us to better explore the effectiveness of AQUA-KV on longer sequences. Since 14 individual tasks are often difficult to analyze, we report the average score across all tasks and provide detailed per-task results in Appendices DE & F. We use the official benchmark code, and evaluate on Instruct models since many LongBench tasks were designed for such models (see Appendix C).
@@ 4.1 Detailed Evaluation & Ablation Analysis
Alternative quantizers. As discussed, AQUA-KV is compatible with any “backbone” quantization scheme. We focus on two schemes described in Section 3.3: HIGGS and Quanto. For HIGGS, we use the quantization group size 1024 and use three grid configurations that have fast GPU support, at 2-, 3- and 4-bit precision (we use $d{=}2$, $n\in\{16,64,256\}$ respectively) and one grid without fast GPU support (2-bit precision, $d{=}4$, $n{=}256$). For Quanto, we use the default group size 64 and per-token compression (0-th axis). Both methods use Round-To-Nearest (RTN) quantization. We evaluate each scheme with and without learned AQUA-KV predictors, and compare against two popular algorithms for KV cache compression: KIVI (Liu et al., 2024c) and KVQuant (Hooper et al., 2024). We also discuss AQUA-KV compatibility with QuaRot and report additional Quanto configurations in Appendix D.
@@
Uncompressed 16 40 6.98 5.61 2.54 7.14 6.13 44.61 48.13 52.92 38.80 46.82
AQUA-KV 2.09 5.7 7.03 5.72 2.62 7.20 6.17 44.30 47.77 52.79 38.31 46.43
HIGGS 2.02 5.1 7.47 5.89 2.77 7.93 8.08 42.80 47.37 52.18 30.92 25.97
@@ Appendix C On LongBench Evaluation on non-Instruct models
Quanto 2-bit gs64 2.50 21.56 21.14 27.03 12.83 64.00 13.05 10.57 81.21 15.54 0.12 22.30 6.01 11.63 5.51 10.84 15.30
AQUA-KV (Quanto gs64) 2.64 10.33 26.24 41.59 11.36 69.00 10.60 21.85 88.33 22.91 0.00 32.67 6.93 12.97 6.65 12.75 29.69
@@
We provide scores on each of the LongBench tasks in Table 5, as well as some additional experiments. The table is supplemented with results on QuaRot quantization method (Ashkboos et al., 2024) that combines Randomized Hadamard Transform (RHT) with standard non-vector quantization. We also report Quanto with larger group size (1024); ablation with post-RoPE quantization and additonal KVSharer setup with 2 shared pairs. The table also contains an ablation on the predictor architecture, including MLP predictors without Layer Nomalization, reduced rank regression (RRR with rank 256), quantized predictors and a corner case with bias-only predictor. We also perform ablation analysis of keeping a longer sequence of unquantized tokens at the beginning of the sentence (4 and 64 tokens). Finally, we evaluate AQUA-KV with different predictor input configurations: removing one of the predictor components or using more distant past layers.
@@
AQUA-KV (Quanto gs64) 2.64 10.33 43.64 41.29 39.99 71.50 53.72 26.15 88.68 23.86 3.53 50.61 24.96 40.38 88.50 24.26 33.54
AQUA-KV (Quanto gs1024) 2.17 53.35 43.66 42.11 39.53 71.00 51.26 25.84 88.67 24.18 4.06 49.36 26.86 39.52 89.50 25.36 33.96
AQUA-KV (QuaRot gs1024) 2.17 8.44 44.54 42.31 39.28 72.00 53.21 25.77 88.48 24.03 4.50 52.26 27.23 40.89 94.00 25.72 33.94
@@
Before RoPE Quanto gs64 2.64 10.90 44.54 42.57 41.31 71.50 53.10 26.02 88.16 24.01 3.53 50.59 26.91 41.50 95.50 25.10 33.69
After RoPE Quanto gs64 2.64 10.33 43.64 41.29 39.99 71.50 53.72 26.15 88.68 23.86 3.53 50.61 24.96 40.38 88.50 24.26 33.54
@@
Uncompressed 16 7.14 38.80 44.45 38.64 68.00 46.60 22.60 87.60 22.94 3.00 49.29 20.53 37.07 49.00 21.88 31.64
AQUA-KV 2.44 7.20 38.31 43.43 37.43 68.50 46.35 22.69 88.17 23.23 3.00 48.93 18.11 36.98 47.50 21.71 30.34
HIGGS 2.06 7.92 30.92 38.73 34.81 48.00 44.54 18.11 84.40 20.26 2.50 38.24 16.17 32.51 15.00 19.34 20.26
@@
Uncompressed 16 6.13 46.82 45.77 46.94 72.00 57.72 23.90 89.42 23.56 8.00 52.58 30.35 43.78 100.00 29.49 31.93
AQUA-KV 2.48 6.17 46.43 45.99 45.71 72.00 56.88 23.84 89.18 23.44 8.50 52.15 29.63 42.77 99.50 29.21 31.15
HIGGS 2.03 8.08 25.97 26.98 17.98 52.50 31.67 12.04 55.46 13.14 7.25 27.19 14.55 22.22 55.79 14.72 12.05
@@
AQUA-KV 4.43 6.14 46.77 45.81 46.87 72.00 57.71 24.06 89.83 23.83 8.00 52.41 30.35 43.37 100.00 28.65 31.86
HIGGS 4.03 6.88 11.54 11.80 7.62 37.75 10.13 4.39 22.54 10.52 7.20 10.50 4.83 5.85 16.26 7.11 5.02
KIVI 4.25 6.14 33.40 46.20 9.61 71.00 10.34 22.27 89.63 21.13 4.58 30.56 6.85 12.79 98.25 12.06 32.39
@@ Appendix F Additional Evaluations of AQUA-KV with H2O
Table 14: Evaluation of Llama 3.2 3B Instruct with H2O pruning mixed with various Key-Value cache compression strategies. A 50% KV cache budget for H2O was used for all evaluations. The left panel contains the average LongBench score for the model. The right panel reports detailed per-task LongBench accuracies and F1 scores for the model.
Config Quant. LongBench Avg $\uparrow$
SamSum
Bits (instruct model)
H2O 16 39.69 42.84 37.87 68.5 44.57 24.84 88.22 22.51 8.5 48.87 16.72 37.22 65 19.35 30.59
H2O + AQUA-KV 2.09 39.11 42.89 38.04 68.5 45.12 24.38 88.64 21.82 6 49.15 16.59 34.65 63.5 19.68 28.62
H2O + HIGGS 2.02 37.42 39.72 36.61 68.5 44.01 24.15 87.52 21.77 5 47.9 16.01 32.32 54.5 18.85 27.07
H2O + AQUA-KV 3.06 39.61 42.59 38.68 68.5 44.58 25.06 88.39 22.17 7.5 49.22 17.22 36.46 64.5 19.42 30.29
H2O + HIGGS 3.02 39.21 42.38 37.62 68 45.04 24.75 89.06 22.32 9 47.96 16.14 36.21 61 19.5 29.93
H2O + AQUA-KV 4.02 39.72 42.65 37.8 68.5 44.58 25.07 89.08 22.35 8.5 49.06 17.06 36.83 64 20.07 30.52
H2O + HIGGS 4.02 39.45 42.23 38.1 68.5 44.67 25 88.53 22.37 9 49.78 16.16 36.85 61 19.81 30.3
Table 15: Evaluation of Llama 3.1 8B Instruct with H2O pruning mixed with various Key-Value cache compression strategies. A 50% KV cache budget for H2O was used for all evaluations. The left panel contains the average LongBench score for the model. The right panel reports detailed per-task LongBench accuracies and F1 scores for the model.
Config Quant. LongBench Avg $\uparrow$
SamSum
Bits (instruct model)
H2O 16 42.35 44.36 44.46 69.5 43.68 25.53 91.83 23.15 6.49 52.93 24.8 43.63 68 22.78 31.79
H2O + AQUA-KV 2.09 42.08 43.87 43.32 69 44.27 25.31 92.38 23.45 6.14 53.2 24.86 41.8 68.5 22.7 30.31
H2O + HIGGS 2.02 41.61 41.69 44.87 69.5 43.55 25.33 91.52 22.47 6.55 52.43 24.08 41.11 68 22.24 29.13
H2O + AQUA-KV 3.05 42.37 44.1 44.65 69.5 44.12 25.38 92.62 23.47 6.5 53.57 24.73 43.3 68 22.13 31.14
H2O + HIGGS 3.02 42.23 43.84 45.28 69.5 43.94 25.34 91.85 23.07 6.49 53.16 24.85 42.01 68 22.41 31.45
H2O + AQUA-KV 4.02 42.37 44.3 44.34 69.5 43.73 25.78 91.72 23.33 6.71 53.08 24.7 43.56 68 22.82 31.64
H2O + HIGGS 4.02 42.17 44.38 43.65 69.5 43.54 25.81 91.93 22.84 6.76 53.25 24.54 42.53 68 22.13 31.58